# ODE: Open-Set Evaluation of Hallucinations in Multimodal Large Language Models

### v3

@@ V-A Hallucination Tendencies in Evaluation Results
V-B ODE in Domain-Specific Scenarios
Our framework also focuses on hallucination detection in specific scenarios or data-scarce fields. The input flexibility of ODE allows for the selection of particular concepts, combined with its ability to generate customizable images, opening new possibilities for tailored data. As multimodal models expand into fields such as autonomous driving and healthcare, constructing diverse, challenging samples becomes critical to overcome existing datasets’ limitations, such as narrow distributions and small sample sizes, which limit adaptability to complex scenarios. Fig. 6 demonstrates our framework’s capability to generate rare concept combinations for the transportation domain, supplementing real-world scene data. This generative capability exposes models to scarce data during training, enhancing real-world adaptability and mitigating overfitting risks. Additionally, fine-tuning with ODE-generated, distributionally diverse images improves model reliability in specialized domains.
@@ V-B ODE in Domain-Specific Scenarios
V-C Fine-Tuning Enhancements
Furthermore, fine-tuning MLLMs with ODE-generated data effectively mitigates object hallucinations. On one hand, the ODE-generated test set not only avoids data contamination while offering rich distributional diversity, uncovering ”thinking patterns” that drive hallucinations in specific scenarios. Based on this, targeted fine-tuning using the erroneous sample set can effectively mitigate the model’s shortcomings. We utilize a dynamic, iteratively generated test set, which helps prevent model overfitting and the potential for ”score boosting” on a single dataset. In our experiments, we extracted the erroneous samples exhibiting hallucinations from the initial round of generated data in the LLaVA model tests and performed fine-tuning on LLaVA. Subsequently, we evaluated the fine-tuned model using the existing AMBER test set. The results presented in Tables IV and V indicate that the fine-tuned model shows significant improvements in performance across both generative and discriminative tasks.
TABLE IV: Generative Task Performance Comparison
@@ V-C Fine-Tuning Enhancements
TABLE V: Discriminative Task Performance Comparison
@@
The dataset, comprising Standard, Long-tail, Random, and Fictional distributions, enabled an open-set analysis. For the object existence hallucination task, 800 sample pairs from each distribution type were used for fine-tuning. The fine-tuned model exhibited better hallucination control Fig. 7 Results indicate that high-frequency distribution samples had the greatest positive effect, while long-tail samples contributed less than random samples. This discrepancy may arise from the regularity of high-frequency data, which facilitates model pattern recognition, whereas the imbalance in long-tail samples can lead to overfitting of frequent categories. Random samples provided better regularization, improving performance across diverse test sets. Incorporating diverse distributions in training data enhances robustness, generalization, and representation of low-frequency objects, boosting overall performance.

### v4

@@ 5.1 Hallucination Tendencies in Evaluation Results
5.2 ODE in Domain-Specific Scenarios
Our framework also addresses hallucination detection in specific scenarios and data-scarce fields. The input flexibility of ODE allows for the selection of particular concepts, combined with its ability to generate customizable images, opening new possibilities for tailored data. As multimodal models expand into fields such as autonomous driving and healthcare, constructing diverse, challenging samples becomes critical to overcome existing datasets’ limitations, such as narrow distributions and small sample sizes, which limit adaptability to complex scenarios. Fig. 6 demonstrates our framework’s capability to generate rare concept combinations for the transportation domain, supplementing real-world scene data. This generative capability exposes models to scarce data during training, thereby enhancing real-world adaptability and mitigating overfitting risks. Additionally, fine-tuning with ODE-generated, distributionally diverse images improves model reliability in specialized domains.
@@ 5.2 ODE in Domain-Specific Scenarios
5.3 Fine-Tuning Enhancements
Furthermore, fine-tuning MLLMs with ODE-generated data effectively mitigates object hallucinations. The ODE-generated test set avoids data contamination while offering rich distributional diversity, uncovering "thinking patterns" that drive hallucinations in specific scenarios. Based on this, targeted fine-tuning using the erroneous sample set can effectively mitigate the model’s shortcomings. We utilize a dynamic, iteratively generated test set, which helps prevent model overfitting and the potential for "score boosting" on a single dataset. In our experiments, we extracted the erroneous samples exhibiting hallucinations from the initial round of generated data in the LLaVA model tests and performed fine-tuning on LLaVA. Subsequently, we evaluated the fine-tuned model using the existing AMBER test set. The results presented in Tables 4 and 5 indicate that the fine-tuned model shows significant improvements in performance across both generative and discriminative tasks.
Table 4: Generative Task Performance Comparison
@@ 5.3 Fine-Tuning Enhancements
Table 5: Discriminative Task Performance Comparison
@@
The dataset, comprising Standard, Long-tail, Random, and Fictional distributions, enabled an open-set analysis. For the object existence hallucination task, 800 sample pairs from each distribution type were used for fine-tuning. The fine-tuned model exhibited better hallucination control. Fig. 7 Results indicate that high-frequency distribution samples had the greatest positive effect, while long-tail samples contributed less than random samples. This discrepancy may arise from the regularity of high-frequency data, which facilitates model pattern recognition, whereas the imbalance in long-tail samples can lead to overfitting of frequent categories. Random samples provided better regularization, improving performance across diverse test sets. Incorporating diverse distributions in training data enhances robustness, generalization, and representation of low-frequency objects, boosting overall performance.
@@ 6 Conclusions
Appendix A Contamination and Synthetic Validity
Fig. 8 highlights the performance of MiniGPT-4 and InstructBLIP on discriminative tasks across COCO 2014, Internet, and ODE-generated datasets. Internet data, as the most recent and reliably crawled dataset, avoids contamination, making it a more trustworthy baseline compared to COCO 2014, which shows artificially inflated results due to potential overlaps with training data. While ODE-generated synthetic data exhibits slightly lower performance than Internet data, feature space validates the close similarity between synthetic and natural images. This confirms that synthetic data can effectively replicate real-world distributions. Moreover, its controllability and distribution diversity establish it as a valuable resource for evaluating model reliability, particularly in mitigating data contamination and enabling the creation of novel and challenging testing scenarios.
Figure 8: Model performances on discriminative tasks.
The degradation results primarily indicate potential data contamination in certain models. Since the Internet and synthetic images do not share entirely identical semantics, some degree of discrepancy is expected due to differences in specific concepts within the images.
Figure 9: Feature Visualization.
We compared the features of three types of images: real images before contamination, contaminated images, and synthetic images generated on the basis of detailed semantic descriptions. The visualization shows that contaminated images form a distinct distribution, while synthetic images align closely with uncontaminated real images. Thus, we conclude that synthetic images are suitable for hallucination evaluation, a perspective also supported by previous works [31].
Appendix B Evaluation Results
Table 6 summarizes the detailed evaluation results of five models under the following conditions: Standard, Random, Fictional, and Long-tail distributions. Tables 7, 8, 9, and 10 present the evaluation results of object-level hallucinations on ODE at the attribute level.
From the detailed results, we derive the following key observations:
  • High Variability Across Attributes: Models excel in state-related attributes due to strong semantic associations in training data. However, action and number attributes expose significant weaknesses, especially in rare or unseen scenarios.
  • Impact of Data Distribution: Standard distributions enable high performance as a result of frequent exposure during training. In contrast, Fictional and Random distributions cause sharp performance declines, exposing over-reliance on memorized correlations. Long-tail distributions further highlight the models’ inability to generalize effectively to sparse data.
  • Task-Specific Limitations: Generative tasks emphasize semantic fluency, whereas discriminative tasks require detailed attribute recognition. These differences result in varying performance gaps between task types.
  • Model-Specific Trends: LLaVA-1.5 demonstrates the best overall performance with balanced precision and recall across attributes and distributions. InstructBLIP achieves high precision but suffers from poor recall, indicating a tendency toward overfitting. MiniGPT-4 and mPLUG-Owl struggle significantly with generalization, particularly in Random and Fictional contexts.
  • Broader Implications for Hallucination Mitigation: Addressing relational hallucination requires improving the diversity and balance of training datasets. Attribute-specific fine-tuning and robust data augmentation strategies are essential for better generalization.
Table 6: Evaluation results of five models (CogVLM, LLaVA-1.5, mPLUG_Owl, MiniGPT-4, and InstructBLIP) under the following conditions: Standard, Random, Fictional, and Long-tail distributions. Each task type—Generative Task (CHAIR, Cover, Hal, Cog) and Discriminative Task (Existence and Attribute)—is evaluated with Accuracy (Acc), Precision (P), Recall (R), and F1-score (F1).
Generative Task Discriminative-Existence Task Discriminative-Attribute Task Criterion Model CHAIR Cover Hal Cog Acc P R F1 Acc P R F1 Standard CogVLM 51.9 76.5 89.1 12.2 50.7 100.0 26.2 41.5 55.2 46.8 55.6 50.8 LLaVA-1.5 38.9 77.7 82.7 8.6 69.5 97.8 55.4 70.7 62.9 32.6 71.9 44.8 mPLUG 50.8 77.2 96.0 11.5 41.7 94.7 13.4 23.5 72.5 41.7 28.6 33.9 MiniGPT-4 49.4 76.0 93.6 14.2 64.5 97.5 48.0 64.3 69.5 39.7 12.5 19.0 InstructBLIP 59.9 75.7 88.1 11.0 66.7 96.8 51.7 67.4 60.8 28.5 51.2 36.6 InternVL-2.5 38.5 60.9 72.3 4.0 75.7 96.1 66.3 78.4 65.9 48.1 87.0 61.9 Qwen2-VL 44.2 75.2 88.6 11.6 75.2 97.0 64.9 77.7 67.0 49.6 86.8 63.1 Cambrian 36.3 64.9 79.2 7.3 76.9 96.5 67.8 79.6 65.7 47.7 90.4 62.4 Random CogVLM 58.1 57.7 87.6 6.0 40.0 89.7 18.0 30.0 57.1 82.5 38.2 52.2 LLaVA-1.5 45.2 57.7 84.2 4.7 74.7 89.7 69.6 78.3 75.7 90.1 65.9 76.1 mPLUG 57.9 56.4 92.1 6.3 40.0 84.0 10.8 19.1 48.4 76.5 20.2 31.9 MiniGPT-4 50.3 57.9 82.2 5.8 66.4 86.9 58.2 69.7 45.0 74.5 10.8 18.8 InstructBLIP 55.9 58.4 83.2 5.6 64.6 87.6 54.6 67.2 72.3 91.5 58.6 71.4 InternVL-2.5 43.0 58.4 67.8 3.7 81.4 89.9 80.7 85.0 72.8 67.0 80.6 73.1 Qwen2-VL 38.5 60.9 72.3 4.0 82.8 88.8 83.5 86.0 71.4 65.8 82.6 73.2 Cambrian 30.9 56.9 61.4 2.3 80.7 87.0 83.0 84.9 71.2 63.4 84.3 72.3 Fictional CogVLM 54.0 55.9 80.7 6.3 39.5 90.9 16.1 27.3 50.5 75.4 28.7 41.5 LLaVA-1.5 48.0 54.2 79.2 4.8 72.5 87.9 66.7 75.8 73.0 88.5 62.0 72.9 mPLUG 59.2 54.0 90.6 6.6 42.5 95.7 11.8 21.0 45.8 72.1 17.4 28.0 MiniGPT-4 48.0 55.4 79.2 7.3 67.1 87.4 58.1 69.8 42.9 75.0 9.7 17.2 InstructBLIP 59.7 56.9 80.2 5.9 66.0 88.6 56.2 68.7 71.0 88.8 58.4 70.4 InternVL-2.5 43.7 55.4 65.3 4.5 78.6 86.5 79.3 82.7 66.2 58.0 73.7 64.9 Qwen2-VL 41.5 60.4 69.8 5.3 77.2 86.6 76.6 81.2 65.8 59.0 75.0 66.0 Cambrian 34.5 55.0 63.4 3.4 79.1 86.6 80.1 83.2 67.1 58.3 77.2 66.4 Long-tail CogVLM 54.8 67.1 89.6 14.3 41.4 90.0 15.6 26.6 61.5 77.3 49.0 59.9 LLaVA-1.5 44.5 71.3 91.1 11.2 51.3 84.7 32.9 47.4 78.2 85.5 74.3 79.5 mPLUG 51.9 71.0 96.0 10.7 38.1 91.4 7.9 14.5 48.9 67.4 24.8 36.2 MiniGPT-4 48.1 75.2 93.1 16.5 63.4 92.9 48.8 63.9 48.0 75.9 17.1 27.9 InstructBLIP 51.3 71.0 87.1 11.2 51.5 91.1 30.4 45.6 73.3 84.6 65.1 73.5 InternVL-2.5 50.4 70.3 86.1 11.1 70.1 88.1 63.9 74 65.2 50.5 89.7 64.6 Qwen2-VL 49.8 72.3 93.1 16.7 69.3 92.2 58.9 71.8 61.4 48.2 82.9 60.9 Cambrian 45.0 53.0 80.7 9.0 68.5 89.6 59.7 71.6 60.8 47.0 88.7 61.4
Table 7: The detailed evaluation results of ODE on the object hallucination (attribute-level) under Standard distribution.
Metric mPLUG-Owl MiniGPT-4 LLaVA-1.5 CogVLM InstructBLIP State Acc 57.8 51.3 80.5 69.5 71.3 P 75.3 84.2 86.8 87.0 86.8 R 28.6 12.5 71.9 55.6 51.2 F1 41.4 21.7 78.6 67.8 64.4 Number Acc 40.4 39.9 92.6 54.5 87.0 P 82.7 93.5 80.2 98.4 98.5 R 15.3 10.6 90.3 32.7 81.9 F1 25.8 19.0 94.1 49.1 89.4 Action Acc 53.3 43.9 66.9 60.1 67.3 P 67.8 45.8 74.2 66.9 74.5 R 17.5 9.6 51.8 51.3 52.6 F1 27.8 15.8 61.0 58.0 61.6
Table 8: The detailed evaluation results of ODE on the object hallucination (attribute-level) under Long-tail distribution.
Metric mPLUG-Owl MiniGPT-4 LLaVA-1.5 CogVLM InstructBLIP State Acc 55.7 50.5 75.5 66.1 64.2 P 69.1 71.4 82.5 84.3 74.6 R 32.2 13.7 64.7 53.4 43.2 F1 43.9 23.0 72.5 65.3 54.7 Number Acc 40.1 42.4 86.8 49.8 83.7 P 78.9 85.7 95.5 91 97.2 R 14.9 16.3 94.2 27.5 78 F1 25 27.44 89.4 42.2 86.5 Action Acc 48.5 51.9 64.4 61 64.4 P 50.7 62.8 64.2 61.8 64.6 R 27.3 20.5 65.2 71.2 63.6 F1 35.4 30.9 64.6 66.1 64.0
Table 9: The detailed evaluation results of ODE on the object hallucination (attribute-level) under Fictional distribution.
Metric mPLUG-Owl MiniGPT-4 LLaVA-1.5 CogVLM InstructBLIP State Acc 51.3 50.2 71.1 55.0 66.4 P 67.5 76.5 87.7 84.4 85.7 R 22.4 11.2 49.1 28.0 41.4 F1 33.6 19.5 62.9 42 55.8 Number Acc 39.9 35.8 76.9 46 73.4 P 85.7 72.7 91.5 86.6 93.6 R 13.4 5.9 72.0 24.0 65.1 F1 23.2 10.9 80.5 37.6 76.7 Action Acc 45.7 46.7 56.5 42.4 57.6 P 30.0 90.9 60. 0 46.5 62.5 R 6.5 21.7 39.1 43.5 43.5 F1 10.7 35.0 47.3 44.9 51.2
Table 10: The detailed evaluation results of ODE on the object hallucination (attribute-level) under Random distribution.
Metric mPLUG-Owl MiniGPT-4 LLaVA-1.5 CogVLM InstructBLIP State Acc 55.0 53.5 71.9 60.3 66.1 P 76.4 78.3 86.3 85.0 87.5 R 22.7 14.9 52.1 39.7 40.5 F1 35.0 25.0 64.9 54.1 55.3 Action Acc 40.9 36.8 80.4 48.0 75.2 P 79.8 78.4 92.5 87.2 95.0 R 16.6 7.2 76.7 27.0 66.8 F1 27.5 13.2 83.8 41.2 78.4 Number Acc 54.7 45.3 54.7 65.6 60.9 P 75 0 75 76.9 76.9 R 18.7 0 18.7 62.5 31.2 F1 29.9 0 29.9 68.9 44.3